OpenStudy (anonymous):

Evaluate (ab)^2 for a=4 and b=3 A. 36 B. 24 C. 81 D. 144

OpenStudy (anonymous):

What is 3 * 4?

OpenStudy (anonymous):

(3*4)^2=12^2=144. So answer is D